OpenStudy (brittany16):

What number needs to be eliminated to solve this equation? -8/X=72

jhonyy9 (jhonyy9):

-8=72x x=-8/72 x=-1/9

jhonyy9 (jhonyy9):

- so but if you divide both sides by 8 you will get this result too

jhonyy9 (jhonyy9):

-8=72x /:8 -1=9x -1/9=x